OpenStudy (anonymous):

The equation below has a solution near x = 0. By replacing the left side of the equation by its linearization, find an approximate value for the solution. Note: Your answer should be in fraction form. e^x+14x+7

OpenStudy (tkhunny):

First of all, we'll need an equation. Otherwise,"solution" has no meaning. Did you mean e^x + 14x + 7 = 0 ? Maybe something else.

OpenStudy (anonymous):

well i guess thats means the same right?

OpenStudy (anonymous):

so i find the linearization of that?

OpenStudy (anonymous):

near x+0? or near =7?

OpenStudy (tkhunny):

No, not at all. Language and terminology are importaant. It enables us to communicate with each other. This is an expression: e^x+14x+7 This is an Equation: e^x+14x+7 = 0 Well, 14x + 7 is already linear, so all we need to focus on is the e^x. How do you find the "linearization" of e^x around x - 0?

OpenStudy (tkhunny):

The problem statement says there is a solution around x = 0.

OpenStudy (anonymous):

1x+1

OpenStudy (anonymous):

right?

OpenStudy (tkhunny):

How did you get that? Is there a derivative in there and your just not sharing your secret?

OpenStudy (anonymous):

f'(a)(x-a)+f(a)

OpenStudy (anonymous):

local linearization formula

OpenStudy (tkhunny):

You have it. Then the linearization of the entire left-hand side is what?

OpenStudy (anonymous):

1x+1?

OpenStudy (tkhunny):

No, that's just e^x. The linearization of e^x is 1+x around x = 0, as we have established. This gives (1+x) + 14x + 7 as the linearization of the entire left-hand side.

OpenStudy (anonymous):

and now what?

OpenStudy (anonymous):

i'm lost

OpenStudy (tkhunny):

The point of linearization is to make the evaluation of the function easier, understanding that we introduce a little error. Unfortunately, we don't quite know the original problem on this one. For x = 0, we have e^0 + 14(0) + 7 = 1 + 0 + 7 = 8 For the linearization, we have (1+0) + 14(0) + 7 = 8 For that, we introduced no error. Now, we need to know the original problem statement in its entirety. What is the original equation?